Grade 12General Physics

A particle subjected to two equal forces along two different directions. If one of the forces is halved, the angle which the resultant makes with th other is also halved. The angle between the forces is-
(a) 45 (b) 60 (c)90 (d) 120
sir is there ay other method other than substitution? Supposing it is a subjective question.
sir pls explain all the steps to me.
thanks a lot
jai

Jai Mahajan
12 Years agoGrade 12

1 Answer

Sumit Majumdar
12 Years ago
Dear student,
?Let us assume that the angle between the given forces be θ. As the forces are equal in magnitude (say F), the resultant R must be making angle θ/2 from each of them. Now let us use triangle law of addition of two vectors. First draw the first force with magnitude 'F' then draw second one F/2. Thus, you would be able to see it clearly that the new resultant R1 will make angle of θ/4 with force 'F'. Please notice that both the forces will still make angle θ between them (only the magnitude of second force has been changed). So, angle inside the triangle between force will be π- θ as earlier. However remaining two angles which were earlier equal to θ/2 have now become θ/4 and 3θ/4 respectively. Now, using the sine law, we get
R_{1}sin\left ( \pi-\theta \right )=\frac{\frac{F}{2}}{sin\left ( \frac{\theta}{4} \right )}=\frac{F}{sin\left ( \frac{3\theta}{4} \right )}
Hence, we get:
2sin\left ( \frac{\theta}{4} \right )=sin\left ( \frac{3\theta}{4} \right )
this helps us in getting the final answer as 120degrees.
Regards
Sumit